# Helical Luttinger liquid on a space-time lattice

### 19 Jan 2024 v1

@@ Abstract
The Luttinger model is a paradigm for the breakdown due to interactions of the Fermi liquid description of one-dimensional massless Dirac fermions. Attempts to discretize the model on a one-dimensional lattice have failed to reproduce the established bosonization results, because of the fermion-doubling obstruction: A local discretization of the Hamiltonian introduces a spurious second species of low-energy excitations, while a nonlocal discretization opens a single-particle gap at the Dirac point. Here we show how to work around this obstruction, by discretizing both space and time to obtain a local action for a helical Luttinger liquid with Hubbard interaction. The approach enables quantum Monte Carlo simulations of interacting relativistic fermions.
@@
Introduction — A quantum spin Hall insulator [1] supports a one-dimensional (1D) helical edge mode of counterpropagating massless electrons (Dirac fermions), with a linear dispersion $E=\pm\hbar v_{\rm F}k$. The crossing at momentum $k=0$ (the Dirac point) is protected from gap-opening by time-reversal symmetry [2] — provided that there is only a single species of low-energy excitations. This topological protection is broken on a lattice by fermion doubling [3]: Any local discretization of the momentum operator $k=-id/dx$ introduces a second Dirac point in the Brillouin zone, an obstruction known as the Nielsen-Ninomiya “no-go” theorem [4].
@@
Fermion doubling is problematic if one wishes to study interaction effects of 1D massless electrons (a Luttinger liquid [5, 6, 7, 8]) by means of a lattice fermion method such as quantum Monte Carlo [9, 10]. One approach is to simulate a 2D system in a ribbon geometry, so that the two fermion species are spatially separated on opposite edges [11, 12, 13, 14]. This is computationally more expensive than a fully 1D simulation, but more fundamentally, the presence of states in the bulk may obscure the intrinsically 1D physics of a Luttinger liquid [15]. A 1D simulation using a nonlocal spatial discretization [16] that avoids fermion doubling was studied recently [17], without success: The nonlocality opens a spurious gap at the Dirac point [17, 18].
@@
Here we show that it can be done: A 1D helical Luttinger liquid can be simulated on a lattice if both space and time are discretized in a way that preserves the locality of the model. The time discretization (in units of $\tau$) allows us to push the second Dirac point up to energies of order $\hbar/\tau$, where it does not affect the low-energy physics — as we demonstrate by comparing the correlators calculated by quantum Monte Carlo with the established results from bosonization [6, 7, 8, 19, 20].
@@
Locally discretized Lagrangian — We construct the space-time lattice using the tangent fermion discretization approach [21, 22, 23, 24, 25], reviewed recently [26]. We first outline that approach for the noninteracting case, in a Lagrangian formulation that is a suitable starting point for the interacting problem.
@@
$C_{\sigma}(x)=Z^{-1}\int{\cal D}\Psi\int{\cal D}\bar{\Psi}\,e^{-{\cal S}[\Psi,\bar{\Psi}]}\bar{\Psi}_{\sigma}(x,0)\Psi_{\sigma}(0,0),$ (11)
@@
irrespective of how space is discretized. This deficiency of the sawtooth (slac) approach was noted in Ref. 17.
@@
Figure 4: Main panels: The data points show the quantum Monte Carlo results for the correlator $R_{x}(x)=\langle\bm{\psi}^{\dagger}(x)\bm{\sigma}_{x}\bm{\psi}(x)\bm{\psi}^{\dagger}(0)\bm{\sigma}_{x}\bm{\psi}(0)\rangle$ of the helical Luttinger liquid, on the space-time lattice with parameters $\beta/\tau=34$, $L/a=71$, $v_{\rm F}=a/\tau$. The different colors refer to different Hubbard interaction strengths $\kappa=Ua/2\pi v_{\rm F}$, repulsive on the left panel and attractive on the right panel. In the latter case the correlator $R_{x}$ changes sign, the plot shows the absolute value on a semi-log scale. The $x$-dependence at $x$ and $L-x$ is the same, because of the periodic boundary conditions, so only the range $20<x<L/$ 2 is plotted. The numerical data on the lattice is compared with the analytical bosonization theory in the continuum (dashed curves [33]). The inset in the left panel combines data for both repulsive and attractive interactions on a log-log scale, to compare with the power law decay (Helical Luttinger liquid on a space-time lattice) (dashed lines).
Conclusion — In summary, we have shown that it is possible to faithfully represent an interacting Luttinger liquid on a lattice, without compromising the fundamental symmetries of massless fermions. The key step is a space-time discretization of the action which is local but does not introduce a spurious second species of massless low-energy fermions. We have tested the validity of this “tangent fermion” approach in the simplest setting without disorder, where we can compare with the known bosonization results in the continuum. We anticipate that tangent fermions can become a powerful tool for the study of interacting topological states of matter.
@@ A.1 Hubbard-Stratonovich transformation of the Euclidean action
To evaluate the fermionic path integral (11) we follow the usual auxiliary field approach [1, 2, 3, 4], by which the two-body Hubbard interaction is transformed into a sum of one-body terms coupled to a fluctuating Ising field $s(x,t)\in\{+1,-1\}$. In a Hamiltonian formulation this is accomplished by the discrete Hubbard-Stratonovich transformation of Ref. 5. We cannot follow that route, because the tangent fermion Hamiltonian is nonlocal, instead we need to work with the Lagrangian formulation — which is local.
@@
$e^{-{\cal S}}=e^{-{\cal S}_{\rm tangent}}\biggl{\langle}\exp\biggl{[}|U\tau|^{1/2}\sum_{x,t}s(x,t)\biggl{(}\bar{\Psi}_{+}(x,t)\Psi_{+}(x,t)-\operatorname{sign}(U\tau)\bar{\Psi}_{-}(x,t)\Psi_{-}(x,t)\biggr{)}\biggr{]}\biggr{\rangle}_{s},$ (23)
where $\langle\cdots\rangle_{s}$ indicates the average over the Ising field $s(x,t)=\pm 1$.
@@
$M_{\sigma}(x,x^{\prime},t,t^{\prime})[s]={}$ $\tfrac{1}{2}\tau^{-1}\bigl{(}-i(1+\cos\hat{k}a)\sin\hat{\omega}\tau+\sigma\gamma(1+\cos\hat{\omega}\tau)\sin\hat{k}a\bigr{)}$
$+\delta_{x,x^{\prime}}\delta_{t,t^{\prime}}\tfrac{1}{16}|U\tau|^{1/2}s(x,t)D^{\dagger}|x,t\rangle\langle x,t|D\times\begin{cases}\sigma&\text{if}\;\;U>0,\\ 1&\text{if}\;\;U<0.\end{cases}$ (26)
@@
$P[s]=\int{\cal D}\Phi\int{\cal D}\bar{\Phi}\,e^{-{\cal S}[\Phi,\bar{\Phi},s]}=\det M_{+}[s]M_{-}[s]$ (27)
for the average over the Ising field $s$. This final average is carried out by means of the Monte Carlo importance sampling algorithm.
@@ B.2 Spin correlator in terms of the bosonic fields
We wish to compute the (equal time) spin correlator
@@
In the second equality we used spin conservation symmetry,
$\begin{split}\langle\psi_{\uparrow}^{\dagger}(x)\psi_{\downarrow}(x)\psi_{\uparrow}^{\dagger}(0)\psi_{\downarrow}(0)\rangle=0,\\ \langle\psi_{\downarrow}^{\dagger}(x)\psi_{\uparrow}(x)\psi_{\downarrow}^{\dagger}(0)\psi_{\uparrow}(0)\rangle=0,\end{split}$ (37)
and translational symmetry,

### 25 Jan 2024 v2

@@ Abstract
The Luttinger model is a paradigm for the breakdown due to interactions of the Fermi liquid description of one-dimensional massless Dirac fermions. Attempts to discretize the model on a one-dimensional lattice have failed to reproduce the established bosonization results, because of the fermion-doubling obstruction: A local and symmetry-preserving discretization of the Hamiltonian introduces a spurious second species of low-energy excitations, while a nonlocal discretization opens a single-particle gap at the Dirac point. Here we show how to work around this obstruction, by discretizing both space and time to obtain a local Lagrangian for a helical Luttinger liquid with Hubbard interaction. The approach enables quantum Monte Carlo simulations that preserve the topological protection of an unpaired Dirac cone.
@@
Introduction — A quantum spin Hall insulator [1] supports a one-dimensional (1D) helical edge mode of counterpropagating massless electrons (Dirac fermions), with a linear dispersion $E=\pm\hbar v_{\rm F}k$. The crossing at momentum $k=0$ (the Dirac point) is protected from gap-opening [2] — provided that there is only a single species of low-energy excitations and provided that fundamental symmetries (time-reversal symmetry, chiral symmetry) are preserved. This topological protection is broken on a lattice by fermion doubling [3]: Any local and symmetry-preserving discretization of the momentum operator $k=-i\hbar d/dx$ must introduce a spurious second Dirac point [4].
@@
Fermion doubling is problematic if one wishes to study interaction effects of 1D massless electrons (a Luttinger liquid [5, 6, 7, 8]) by means of a lattice fermion method such as quantum Monte Carlo [9, 10, 11, 12, 13]. One way to preserve the time-reversal and chiral symmetries on a lattice is to simulate a 2D system in a ribbon geometry, so that the two fermion species are spatially separated on opposite edges [14, 15, 16, 17, 18]. The 2D simulation is computationally more expensive than a fully 1D simulation, but more fundamentally, the presence of states in the bulk may obscure the intrinsically 1D physics of a Luttinger liquid [19]. A 1D simulation using a nonlocal spatial discretization [20] that avoids fermion doubling was studied recently [21], without success: The nonlocality gaps the Dirac point [21, 22].
@@
Here we show that it can be done: A 1D helical Luttinger liquid can be simulated on a lattice if both space and time are discretized in a way that preserves the locality of the Lagrangian. The time discretization (in units of $\tau$) pushes the second Dirac point up to energies of order $\hbar/\tau$, where it does not affect the low-energy physics — as we demonstrate by comparing quantum Monte Carlo simulations with results from bosonization [6, 7, 8, 23, 24].
@@
Locally discretized Lagrangian — We construct the space-time lattice using the tangent fermion discretization approach [25, 26, 27, 28, 29, 30]. We first outline that approach for the noninteracting case, in a Lagrangian formulation that is a suitable starting point for the interacting problem.
@@
$C_{\sigma}(x)=Z^{-1}\int{\cal D}\bar{\Psi}\int{\cal D}\Psi\,e^{-{\cal S}[\Psi,\bar{\Psi}]}\bar{\Psi}_{\sigma}(x,0)\Psi_{\sigma}(0,0),$ (11)
@@
irrespective of how space is discretized. This deficiency of the sawtooth (slac) approach was noted in Ref. 21.
@@
Figure 4: Main panels: The data points show the quantum Monte Carlo results for the correlator $R_{x}(x)=\langle\bm{\psi}^{\dagger}(x)\bm{\sigma}_{x}\bm{\psi}(x)\bm{\psi}^{\dagger}(0)\bm{\sigma}_{x}\bm{\psi}(0)\rangle$ of the helical Luttinger liquid, on the space-time lattice with parameters $\beta/\tau=34$, $L/a=71$, $v_{\rm F}=a/\tau$. The different colors refer to different Hubbard interaction strengths $\kappa=Ua/2\pi v_{\rm F}$, repulsive on the left panel and attractive on the right panel. In the latter case the correlator $R_{x}$ changes sign, the plot shows the absolute value on a log-linear scale. The $x$-dependence at $x$ and $L-x$ is the same, because of the periodic boundary conditions, so only the range $20<x<L/$ 2 is plotted. The numerical data on the lattice is compared with the analytical bosonization theory in the continuum (dashed curves [37]). The inset in the left panel combines data for both repulsive and attractive interactions on a log-log scale, to compare with the power law decay (Helical Luttinger liquid on a space-time lattice) (dashed lines).
@@
Conclusion — We have shown that it is possible to faithfully represent an interacting Luttinger liquid on a lattice, without compromising the fundamental symmetries of massless fermions. The key step is a space-time discretization of the Lagrangian which is local but does not introduce a spurious second species of low-energy excitations. We have tested the validity of this “tangent fermion” approach in the simplest setting where we can compare with the known bosonization results in the continuum. We anticipate that tangent fermions can become a powerful tool for the study of interacting topological states of matter, where it is essential to maintain the topological protection of an unpaired Dirac cone.
@@ A.1 Hubbard-Stratonovich transformation of the Euclidean action
To evaluate the fermionic path integral representation of the partition function,
$Z=\int{\cal D}\Psi\int{\cal D}\bar{\Psi}\,e^{-{\cal S}[\Psi,\bar{\Psi}]}$ (21)
we follow the usual auxiliary field approach [9, 10, 11, 12, 13], by which the two-body Hubbard interaction is transformed into a sum of one-body terms coupled to a fluctuating Ising field $s(x,t)\in\{+1,-1\}$. In a Hamiltonian formulation this is accomplished by the discrete Hubbard-Stratonovich transformation of Ref. 10. We cannot follow that route, because the tangent fermion Hamiltonian is nonlocal, instead we need to work with the Lagrangian formulation — which is local.
@@
$e^{-{\cal S}}=e^{-{\cal S}_{\rm tangent}}\frac{1}{2^{\frac{\beta L}{\tau a}}}\sum_{s(x,t)=\pm 1}\exp\biggl{[}|U\tau|^{1/2}\sum_{x,t}s(x,t)\biggl{(}\bar{\Psi}_{+}(x,t)\Psi_{+}(x,t)-\operatorname{sign}(U\tau)\bar{\Psi}_{-}(x,t)\Psi_{-}(x,t)\biggr{)}\biggr{]},$ (24)
@@
The Jacobian $J=\det\hat{D}^{\dagger}\hat{D}$ of the transformation is independent of the Ising field.
@@
$M_{\sigma}(x,x^{\prime},t,t^{\prime})[s]={}$ $\tfrac{1}{2}\bigl{(}-i(1+\cos\hat{k}a)\sin\hat{\omega}\tau+\sigma\gamma(1+\cos\hat{\omega}\tau)\sin\hat{k}a\bigr{)}$
$+\delta_{x,x^{\prime}}\delta_{t,t^{\prime}}|U\tau|^{1/2}s(x,t)D^{\dagger}|x,t\rangle\langle x,t|D\times\begin{cases}\sigma&\text{if}\;\;U>0,\\ 1&\text{if}\;\;U<0.\end{cases}$ (27)
@@
$P[s]=\int{\cal D}\bar{\Phi}\int{\cal D}\Phi\,e^{-{\cal S}[\Phi,\bar{\Phi},s]}=\det M_{+}[s]M_{-}[s]$ (28)
for the average over the Ising field. This final average is carried out by means of the Monte Carlo importance sampling algorithm.
@@ A.2 Absence of a sign problem
A.3 Correlators
We apply the quantum Monte Carlo calculation to equal-time correlators of the form
$\langle\psi_{\sigma}^{\dagger}(x_{1})\psi_{\sigma}(x_{2})\psi_{\sigma^{\prime}}^{\dagger}(x_{3})\psi_{\sigma^{\prime}}(x_{4})\rangle=\langle(\hat{D}\phi_{\sigma})^{\dagger}(x_{1})(\hat{D}\phi_{\sigma})(x_{2})(\hat{D}\phi_{\sigma^{\prime}})^{\dagger}(x_{3})(\hat{D}\phi)_{\sigma^{\prime}}(x_{4})\rangle$
$\qquad=\int{\cal D}\bar{\Phi}\int{\cal D}\Phi\,\langle e^{-S[\Phi,\bar{\Phi},s]}\rangle_{s}(\bar{\Phi}_{\sigma}\hat{D}^{\dagger})(x_{1})(\hat{D}\Phi_{\sigma})(x_{2})(\bar{\Phi}_{\sigma^{\prime}}\hat{D}^{\dagger})(x_{3})(\hat{D}\Phi)_{\sigma^{\prime}}(x_{4})$
$\qquad=\frac{1}{Z}\sum_{s(x,t)=\pm 1}P[s]\bigg{(}\hat{A}_{\sigma}(x_{2},x_{1})[s]\hat{A}_{\sigma^{\prime}}(x_{4},x_{3})[s]-\delta_{\sigma\sigma^{\prime}}\hat{A}_{\sigma}(x_{4},x_{1})[s]\hat{A}_{\sigma}(x_{2},x_{3})[s]\bigg{)}.$ (31)
In the final equality we defined $\hat{A}_{\sigma}[s]=\hat{D}M_{\sigma}^{-1}[s]\hat{D}^{\dagger}$ and we have used the integration formula (Wick’s theorem)
$\int{\cal D}\bar{\Phi}\int{\cal D}\Phi\,e^{-\bar{\Phi}M\Phi}\Phi_{k}\Phi_{l}\bar{\Phi}_{m}\bar{\Phi}_{n}=(\det M)\bigl{[}(M^{-1})_{kn}(M^{-1})_{lm}-(M^{-1})_{km}(M^{-1})_{ln}\bigr{]}.$ (32)
We consider the spin correlator
[…]
$+\langle\psi_{\downarrow}^{\dagger}(x)\psi_{\downarrow}(0)\psi_{\uparrow}^{\dagger}(0)\psi_{\uparrow}(x)\rangle\big{)}.$ (33)
In the second equality we used spin conservation symmetry,
$\begin{split}\langle\psi_{\uparrow}^{\dagger}(x)\psi_{\downarrow}(x)\psi_{\uparrow}^{\dagger}(0)\psi_{\downarrow}(0)\rangle=0,\\ \langle\psi_{\downarrow}^{\dagger}(x)\psi_{\uparrow}(x)\psi_{\downarrow}^{\dagger}(0)\psi_{\uparrow}(0)\rangle=0.\end{split}$ (34)
We substitute Eq. (31) into to Eq. (33),
$R_{x}(x)=-\tfrac{1}{4}\big{\langle}\hat{A}_{\uparrow}(0,x)[s]\hat{A}_{\downarrow}(x,0)[s]+\hat{A}_{\downarrow}(0,x)[s]\hat{A}_{\uparrow}(x,0)[s]\bigr{\rangle}_{s}.$ (35)
Using the symmetries (29) and (30) of the $M$-matrix we simplify it to
$R_{x}(x)=\begin{cases}\tfrac{1}{4}\bigl{\langle}\big{|}A_{\uparrow}(x,0)[s]\big{|}^{2}+\big{|}A_{\uparrow}(0,x)[s]\big{|}^{2}\bigr{\rangle}_{s}&\text{if }U>0,\\ -\tfrac{1}{2}\operatorname{Re}\bigl{\langle}\{A_{\uparrow}(x,0)[s]A^{*}_{\uparrow}(0,x)[s]\}\bigr{\rangle}_{s}&\text{if }U<0,\end{cases}$ (36)
[…]
The average in Eq. (36) is over the Ising field $s$. To improve the statistics we make use of translational invariance in space and imaginary time, by additionally averaging the correlator over the initial position (replacing $(x,0)\mapsto(x+y,y)$ with $0<y<L$), as well as over $0<t<\beta$.
A.4 Monte Carlo averaging
For the Monte Carlo averaging we perform local updates of the auxiliary Ising field, one spin-flip at the time. An operator average is then sampled at each Monte Carlo iteration, which includes $\beta L/\tau a$ local spin-flip steps.
We have tried out three alternative numerical methods:
  1. 1.
    The first method is a simple dense matrix calculation, where we recalculate $P[s]$ after each spin-flip and $A_{\uparrow}[s]$ after each Monte Carlo step, using neither the sparsity of the $M$-matrix nor the locality of the update. This is optimal for systems of small sizes $(L/a\lesssim 11,\;\beta/\tau\lesssim 6)$.
  2. 2.
    In the second method we use the locality of the update of the Ising field, by employing the Woodbury formula for the update of $P[s]$ and $A_{\uparrow}[s]$. This is favorable for systems of medium sizes $(L/a\approx 41,\,\beta/\tau\approx 20)$.
  3. 3.
    In the third method we use the sparsity of the $M$-matrix, with the help of the SuperLU library. [^1] X. S. Li, An overview of SuperLU: Algorithms, implementation, and user interface, ACM Transactions on Mathematical Software, 31, 302 (2005). This gives the best performance for large systems $(L/a\gtrsim 61,\,\beta/\tau\gtrsim 30)$.
@@ B.2 Spin correlator in terms of the bosonic fields
We wish to compute the (equal-time) spin correlator
@@
In the second equality we used spin conservation symmetry (34) and translational symmetry,